OpenStudy (anonymous):

f(x)=(x+6)^2(x-2)^2 what is the y-intercept of f? Power function of f? Turning points? Behavior of xintercept (2) " " -6

OpenStudy (anonymous):

y-intercept (0,f(0)) = (0,144)

OpenStudy (anonymous):

The degree of f as a polynomial is 4

OpenStudy (anonymous):

\[ f'(x)=4 x^3+24 x^2-16 x-96\\ f''(x)=12 x^2+48 x-16 f''(x) = 0\\ \begin{array}{c} x=\frac{2}{3} \left(-3-2 \sqrt{3}\right) \\ x=\frac{2}{3} \left(-3+2 \sqrt{3}\right) \\ \end{array} \] These are the x's where f has inflection (turning point)

OpenStudy (anonymous):

The function is tangent to the x-axis at x=-6 and x=2 because these are double roots.

OpenStudy (anonymous):

So my behavior of the x-intercept for 2 ?

OpenStudy (anonymous):

The graph is tangent to the x-axis.

OpenStudy (anonymous):

See the attached graph of your function.

OpenStudy (anonymous):

ok, I see so for my behavior of the x-intercept of 2 is y=-6 and 2? and my maxium number of turning points

OpenStudy (anonymous):

Are turning points inflection points?

OpenStudy (anonymous):

f has a local minimum at x =-6 and x =2 it has a local maximum at x=-2

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

I think... so for the x-intercept of (2) It the -6 and 2 for x-intercept(-6) I put -2

OpenStudy (anonymous):

x-intercepts are at x=-6 and x =2

OpenStudy (anonymous):

I have to go bye.

OpenStudy (anonymous):

see my problem says what s he behavior of grahp near the xintercept of 2